Info: There is also a provision in the contract that SM would receive a discount (similar...

Info:

There is also a provision in the contract that SM would receive a discount (similar to that which would be reflected in a separate financing transaction between EYE SPY and SM) from the contract price of $10 million if they pay the cash component within three days of when the contract is signed. EYE SPY determined a discount of $500,000 for this financing based on applying the typical credit rate for the equipment and integration services to be delivered at the end of year one and the monthly delivery of maintenance services in year two through six of the contract.

Question:

How do I calculate the implicit interest rate of the $500,000 discount from $10 million to $9.5 million?
Answer #1

As per the given question contract period is Six years.

If the contract price is paid within three days, price will be $ 95,00,000 and

if it is paid after 6 years price will be $ 1,00,00,000 (assumed it will be paid at the end of the contract period, as no info is given for installments)

Let us assume the implicit interest rate as "X"

At implicit rate of interest, Cash price = Present value of future outflows

9500000 = 10000000 / (1+X)6

(1+X)6 = 10000000 / 9500000

(1+X) = (10000000 / 9500000 )1/6

1+X = 1.00858553

X = 1.00858553 - 1

X = 0.00858553

X = 0.858553 %

Hence, implicit interest rate of the $500,000 discount from $10 million to $9.5 million is 0.858553 %